# Large Language Models for Cyber Security: A Systematic Literature Review

### v3

@@ Abstract.
The rapid advancement of Large Language Models (LLMs) has opened up new opportunities for leveraging artificial intelligence in a variety of application domains, including cybersecurity. As the volume and sophistication of cyber threats continue to grow, there is an increasing need for intelligent systems that can automatically detect vulnerabilities, analyze malware, and respond to attacks. In this survey, we conduct a comprehensive review of the literature on the application of LLMs in cybersecurity (LLM4Security). By comprehensively collecting over 30K relevant papers and systematically analyzing 127 papers from top security and software engineering venues, we aim to provide a holistic view of how LLMs are being used to solve diverse problems across the cybersecurity domain.
Through our analysis, we identify several key findings. First, we observe that LLMs are being applied to a wide range of cybersecurity tasks, including vulnerability detection, malware analysis, network intrusion detection, and phishing detection. Second, we find that the datasets used for training and evaluating LLMs in these tasks are often limited in size and diversity, highlighting the need for more comprehensive and representative datasets. Third, we identify several promising techniques for adapting LLMs to specific cybersecurity domains, such as fine-tuning, transfer learning, and domain-specific pre-training. Finally, we discuss the main challenges and opportunities for future research in LLM4Security, including the need for more interpretable and explainable models, the importance of addressing data privacy and security concerns, and the potential for leveraging LLMs for proactive defense and threat hunting.
@@ 1. Introduction
In particular, one important domain where LLMs are beginning to show promise is cybersecurity. With the growing volume and sophistication of cyber threats, there is an urgent need for intelligent systems that can automatically detect vulnerabilities, analyze malware, and respond to attacks (Chen et al., 2023c; Bilot et al., 2023; Montes et al., 2021). Recent research has explored the application of LLMs across a wide range of cybersecurity tasks, i.e., LLM4Security hereafter. In the domain of software security, LLMs have been used for detecting vulnerabilities from natural language descriptions and source code, as well as generating security-related code, such as patches and exploits. These models have shown high accuracy in identifying vulnerable code snippets and generating effective patches for common types of vulnerabilities (Chow et al., 2023; Fu et al., 2022; Charalambous et al., 2023). Beyond code-level analysis, LLMs have also been applied to understand and analyze higher-level security artifacts, such as security policies and privacy policies, helping to classify documents and detect potential violations (Hartvigsen et al., 2022; Mets et al., 2023). In the realm of network security, LLMs have demonstrated the ability to detect and classify various types of attacks from network traffic data, including DDoS attacks, port scanning, and botnet traffic (Alkhatib et al., 2022; Moskal et al., 2023; Ali and Kostakos, 2023). Malware analysis is another key area where LLMs are showing promise, with models being used to classify malware families based on textual analysis reports and behavioral descriptions, as well as detecting malicious domains and URLs (Joyce et al., 2023; Liu et al., 2023e). LLMs have also been employed in the field of social engineering to detect and defend against phishing attacks by analyzing email contents and identifying deceptive language patterns (Jamal and Wimmer, 2023; Ranade et al., 2021). Moreover, researchers are exploring the use of LLMs to enhance the robustness and resilience of security systems themselves, by generating adversarial examples for testing the robustness of security classifiers and simulating realistic attack scenarios for training and evaluation purposes (Temara, 2023; Charan et al., 2023; Saha et al., 2023). These diverse applications demonstrate the significant potential of LLMs to improve the efficiency and effectiveness of cybersecurity practices by processing and extracting insights from large amounts of unstructured text, learning patterns from vast datasets, and generating relevant examples for testing and training purposes.
@@
Reference Year Scope of topics Dimensions of discourse Time frame Papers
Motlagh et al. (Hou et al., 2024) 2024 Security application Task 2022-2023 Not specified
Divakaran et al (Divakaran and Peddinti, 2024) 2024 Security application Task 2020-2024 Not specified
Yao et al. (Yao et al., 2024) 2023
Security application
Security of LLM
Model
Task
2019-2024 281
Yigit et al. (Yigit et al., 2024) 2024
Security application
Security of LLM
Task 2020-2024 Not specified
Coelho et al. (de Jesus Coelho da Silva and Westphall, 2024) 2024 Security application
Task
Domain specific technique
2021-2023 19
Novelli et al. (Novelli et al., 2024) 2024
Security application
Security of LLM
Task 2020-2024 Not specified
LLM4Security 2024 Security application
Model
Task
Domain specific technique
Data
2020-2024 127
@@
To address these limitations and provide an in-depth understanding of the state-of-the-art in LLM4Security, we conduct a systematic and extensive survey of the literature. By comprehensively collecting 38,112 relevant papers and systematically analyzing 127 papers from top security and software engineering venues, our survey aims to provide a holistic view of how LLMs are being applied to solve diverse problems across the cybersecurity domain. In addition to identifying the types of tasks that LLMs are being used for, we also examine the specific datasets, preprocessing techniques, and domain adaptation methods employed in each case. This enables us to provide a more nuanced analysis of the strengths and limitations of different approaches, and to identify the most promising directions for future research. Specifically, we focus on answering four key research questions (RQs):
@@
  • We conduct a comprehensive Systematic Literature Review (SLR) to investigate the latest research on LLM4Security, providing a mapping of the current landscape. Our search covers an extensive number of over 38,112 papers. With further quality-based and relevance-based filtering, we retain 127 papers for later detailed review.
@@
  • We analyze the distribution of the 127 selected papers across venues and over time, revealing rapid growth in LLM4Security research especially in 2022-2023, and categorizes the characteristics of mainstream LLMs employed in the security domain.
@@ 2.2. Search Strategy
Step2: Establish QGS. In this step, we start with creating a manually curated set of studies that have been carefully screened to form the QGS. A total of 41 papers relevant to LLM4Sec are manually identified, aligning with the research objective and encompassing various techniques, application domains, and evaluation methods.
@@
Step4: Conduct an automated search. These identified keywords are paired one by one and input into automated searches across the above-mentioned seven widely used databases. Our automated search focused on papers published after 2019, in which GPT-2 was published, as it marked a significant milestone in the development of large language models. The search was conducted in the title, abstract, and keyword fields of the papers in each database. Specifically, the number of papers retrieved from each database after applying the search query and the year filter (2019-2023) is as follows: 3,398 papers in ACM Digital Library, 2,112 papers in IEEE Xplore, 724 papers in Science Direct, 4,245 papers in Web of Science, 23,721 papers in Springer, 7,154 papers in Wiley, and 3,557 papers in arXiv.
@@ 2.3. Study Selection
Ex#3: Papers not published in identified conferences or journals, nor as preprints on arXiv; Ex#4: Papers that do not focus on security tasks (e.g., natural language processing tasks in general domains);
@@
After obtaining the initial pool of 38,112 papers (38,071 from the automated search and 41 from the QGS), we conducted a multi-stage study selection process to identify the most relevant and high-quality papers for our systematic review.
@@ 2.3.1. Coarse-Grained Inclusion and Exclusion Criteria
To select relevant papers for our research questions, we defined four inclusion criteria and eight exclusion criteria (as listed in Table 2) for the coarse-grained paper selection process. Among them, In#1, Ex#1, Ex#2, and Ex#3 were automatically filtered based on the keywords, duplication status, length, and publication venue of the papers. The remaining inclusion criteria (In#2~4) and exclusion criteria (Ex#4~8) were manually applied by inspecting the topic and content of each paper. Specifically, the criteria of In#1 retained 7,582 papers whose titles and abstracts contained a pair of the identified search keywords. Subsequently, Ex#1 filtered out 440 duplicate or multi-version papers from the same authors with little difference. Next, the automated filtering criteria Ex#2 was applied to exclude short papers, tool demos, keynotes, editorials, books, theses, workshop papers, or poster papers, resulting in 4,855 papers being removed. The remaining papers were then screened based on the criteria Ex#3, which retained 523 full research papers published in the identified venues or as preprints on arXiv. The remaining inclusion and exclusion criteria (In#2~4, Ex#4~8) were then manually applied to the titles and abstracts of these 523 papers, in order to determine their relevance to the research topic. Three researchers independently applied the inclusion and exclusion criteria to the titles and abstracts. Disagreements were resolved through discussion and consensus. After this manual inspection stage, 156 papers were included for further fine-grained full-text quality assessment.
@@ 2.3.2. Fine-grained Quality Assessment
Each criterion was scored on a 3-point scale (0: not met, 1: partially met, 2: fully met). Papers with a total score of 6 or higher (out of 10) were considered as having acceptable quality. After the quality assessment, 93 papers remained in the selected set.
@@ 2.3.3. Forward and Backward Snowballing
To further expand the coverage of relevant literature, we performed forward and backward snowballing on the 93 selected papers. Forward snowballing identified papers that cited the selected papers, while backward snowballing identified papers that were referenced by the selected papers.
Here we obtained 2,056 and 5,255 papers separately during the forward and backward process. Then we applied the same inclusion/exclusion criteria and quality assessment to the papers found through snowballing. After the initial keyword filtering and deduplication, there were 1,978 papers that remained available. Among them, 68 papers were excluded during the page number filtering step, and 1,235 papers were deleted to ensure the papers were published in the selected venues. After confirming the paper topics and assessing the paper quality, only 44 papers were ultimately retained in the snowballing process, resulting in a final set of 127 papers for data extraction and synthesis.
@@ 2.4. Statistics of Selected Papers
After conducting searches and snowballing, a total of 127 relevant research papers were ultimately obtained. The distribution of the included documents is outlined in Figure 3. As depicted in Figure 3(A), 39% of the papers underwent peer review before publication. Among these venues, ICSE had the highest frequency, contributing 7%. Other venues making significant contributions included FSE, ISSTA, ASE, and TSE, with contributions of 5%, 5%, 3%, and 3% respectively. Meanwhile, the remaining 61% of the papers were published on arXiv, an open-access platform serving as a repository for scholarly articles. This discovery is unsurprising given the rapid emergence of new LLM4Security studies, with many works recently completed and potentially undergoing peer review. Despite lacking peer review, we conducted rigorous quality assessments on all collected papers to ensure the integrity of our investigation results. This approach enables us to include all high-quality and relevant publications while upholding stringent research standards.
@@
Figure 3. Overview of the selected 127 papers’ distribution.
@@ 4 The types and features of datasets used
The temporal distribution of the included papers is depicted in Figure 3(B). Since 2020, there has been a notable upward trend in the number of publications. In 2020, only 1 relevant paper was published, followed by 2 in 2021. However, the number of papers sharply increased to 11 by 2022. Surprisingly, in 2023, the total count surged to 109 published papers. This rapid growth trend signifies an increasing interest in LLM4Security research. Currently, many works from 2024 are still under review or unpublished. Hence, we have chosen only 6 representative papers. We will continue to observe the developments in LLM4Security research throughout 2024.
@@ 3. RQ1: What types of security tasks have been facilitated by LLM-based approaches?
This section delves into the detailed examination of LLM utilization across diverse security domains. We have classified them into six primary domains, aligning with the themes of the collected papers: software and system security, network security, information and content security, hardware security, and blockchain security, totaling 127 papers. Figure 4 visually depicts the distribution of LLMs within these six domains. Additionally, Table 4 offers a comprehensive breakdown of research detailing specific security tasks addressed through LLM application.
@@
The majority of research activity in the realm of software and system security, constituting around 59.84% of the total research output, is attributed to the advancements made by code LLMs (Sadik et al., 2023; Zheng et al., 2024a; Zhang et al., 2024) and the extensive applications of LLMs in software engineering (Hou et al., 2024). This emphasis underscores the significant role and impact of LLMs in software and system security, indicating a predominant focus on leveraging LLMs to automate the handling of potential security issues in programs and systems. Approximately 17.32% of the research focus pertains to network security tasks, highlighting the importance of LLMs in aiding traffic detection and network threat analysis. Information and content security activities represent around 14.17% of the research output, signaling a growing interest in employing LLMs for generating and detecting fake content. Conversely, activities in hardware security and blockchain security account for approximately 4.72% and 3.94% of the research output, respectively, suggesting that while exploration in these domains has been comparatively limited thus far, there remains research potential in utilizing LLMs to analyze hardware-level vulnerabilities and potential security risks in blockchain technology.
@@
Network Security Web fuzzing (3) 22
@@
Software and System Security Vulnerability detection (17) 76
@@
Information and Content Security Phishing and scam detection (8) 18
@@
Hardware Security Hardware vulnerability detection (2) 6
@@
Blockchain Security Smart contract security (4) 5

### v4

@@ Abstract.
The rapid advancement of Large Language Models (LLMs) has opened up new opportunities for leveraging artificial intelligence in a variety of application domains, including cybersecurity. As the volume and sophistication of cyber threats continue to grow, there is an increasing need for intelligent systems that can automatically detect vulnerabilities, analyze malware, and respond to attacks. In this survey, we conduct a comprehensive review of the literature on the application of LLMs in cybersecurity (LLM4Security). By comprehensively collecting over 40K relevant papers and systematically analyzing 185 papers from top security and software engineering venues, we aim to provide a holistic view of how LLMs are being used to solve diverse problems across the cybersecurity domain.
Through our analysis, we identify several key findings. First, we observe that LLMs are being applied to an expanding range of cybersecurity tasks, including vulnerability detection, malware analysis, network intrusion detection, and phishing detection. Second, we analyze how different LLM architectures (such as encoder-only, encoder-decoder, and decoder-only) are utilized across various security domains and over time, identifying key application trends in LLM4Security. Third, we identify several increasingly sophisticated techniques for adapting LLMs to specific cybersecurity domains, such as advanced fine-tuning, prompt engineering, and external augmentation strategies. Fourth, we find that the datasets used for training and evaluating LLMs in these tasks are often limited in size and diversity, highlighting the need for more comprehensive datasets and the emerging use of LLMs for data augmentation. Finally, we discuss the main challenges and opportunities for future research in LLM4Security, including the need for more interpretable and robust models, the importance of addressing data security and the inherent security risks of LLMs themselves, and the potential for leveraging LLMs for proactive defense and threat hunting.
@@ 1. Introduction
In particular, one important domain where LLMs are beginning to show promise is cybersecurity. With the growing volume and sophistication of cyber threats, there is an urgent need for intelligent systems that can automatically detect vulnerabilities, analyze malware, and respond to attacks (Chen et al., 2023c; Bilot et al., 2023; Montes et al., 2021). Recent research has explored the application of LLMs across a wide range of cybersecurity tasks, i.e., LLM4Security hereafter. In the domain of software security, LLMs have been used for detecting vulnerabilities from natural language descriptions and source code, as well as generating security-related code, such as patches and exploits. These models have shown high accuracy in identifying vulnerable code snippets and generating effective patches for common types of vulnerabilities (Chow et al., 2023; Fu et al., 2022; Charalambous et al., 2023). Beyond code-level analysis, LLMs have also been applied to understand and analyze higher-level security artifacts, such as security policies and privacy policies, helping to classify documents and detect potential violations (Hartvigsen et al., 2022; Mets et al., 2023). In the realm of network security, LLMs have demonstrated the ability to detect and classify various types of attacks from network traffic data, including DDoS attacks, port scanning, and botnet traffic (Alkhatib et al., 2022; Moskal et al., 2023; Ali and Kostakos, 2023). Malware analysis is another key area where LLMs are showing promise, with models being used to classify malware families based on textual analysis reports and behavioral descriptions, as well as detecting malicious domains and URLs (Joyce et al., 2023; Liu et al., 2023e). LLMs have also been employed in the field of social engineering to detect and defend against phishing attacks by analyzing email contents and identifying deceptive language patterns (Jamal et al., 2024; Ranade et al., 2021). Moreover, researchers are exploring the use of LLMs to enhance the robustness and resilience of security systems themselves, by generating adversarial examples for testing the robustness of security classifiers and simulating realistic attack scenarios for training and evaluation purposes (Temara, 2023; Charan et al., 2023; Saha et al., 2024). These diverse applications demonstrate the significant potential of LLMs to improve the efficiency and effectiveness of cybersecurity practices by processing and extracting insights from large amounts of unstructured text, learning patterns from vast datasets, and generating relevant examples for testing and training purposes.
@@
Reference Year Scope of topics Dimensions of discourse Papers
Motlagh et al. (Hou et al., 2024) 2024 Security application Task Not specified
Divakaran et al (Divakaran and Peddinti, 2024) 2024 Security application Task Not specified
Yao et al. (Yao et al., 2024) 2024 Security application Security of LLM Model Task 281
Yigit et al. (Yigit et al., 2024) 2024 Security application Security of LLM Task Not specified
Coelho et al. (de Jesus Coelho da Silva and Westphall, 2024) 2024 Security application Task Domain specific technique 19
Novelli et al. (Novelli et al., 2024) 2024 Security application Security of LLM Task Not specified
LLM4Security 2024 Security application Model Task Domain specific technique Data 185
@@
To address these limitations and provide an in-depth understanding of the state-of-the-art in LLM4Security, we conduct a systematic and extensive survey of the literature. By comprehensively collecting 47,135 relevant papers and systematically analyzing 185 papers from top security and software engineering venues, our survey aims to provide a holistic view of how LLMs are being applied to solve diverse problems across the cybersecurity domain. In addition to identifying the types of tasks that LLMs are being used for, we also examine the specific datasets, preprocessing techniques, and domain adaptation methods employed in each case. This enables us to provide a more nuanced analysis of the strengths and limitations of different approaches, and to identify the most promising directions for future research. Specifically, we focus on answering four key research questions (RQs):
@@
  • We conduct a comprehensive Systematic Literature Review (SLR) to investigate the latest research on LLM4Security, providing a mapping of the current landscape. Our search covers an extensive number of over 47,135 papers. With further quality-based and relevance-based filtering, we retain 185 papers for later detailed review.
@@
  • We analyze the distribution of the 185 selected papers across venues and over time, and categorizes the characteristics of mainstream LLMs employed in the security domain.
@@ 2.2. Search Strategy
Step2: Establish QGS. In this step, we start by creating a manually curated set of studies that have been carefully screened to form the QGS. A total of 118 papers relevant to LLM4Sec are manually identified, aligning with the research objective and encompassing various techniques, application domains, and evaluation methods.
@@
Step4: Conduct an automated search. These identified keywords are paired one by one and input into automated searches across the above-mentioned seven widely used databases. Our automated search focused on papers published after 2019, in which GPT-2 was published, as it marked a significant milestone in the development of large language models. The search was conducted in the title, abstract, and keyword fields of the papers in each database. Specifically, the number of papers retrieved from each database after applying the search query and the year filter (2019-2025) is as follows: 5,595 papers in ACM Digital Library, 3,745 papers in IEEE Xplore, 1,242 papers in Science Direct, 5,516 papers in Web of Science, 23,986 papers in Springer, 705 papers in Wiley, and 6,346 papers in arXiv.
@@ 2.3. Study Selection
Ex#3: Papers not published in identified conferences or journals, nor as preprints on arXiv;
Ex#4: Papers that do not focus on security tasks (e.g., natural language processing tasks in general domains);
@@
After obtaining the initial pool of 47,135 papers (47,017 from the automated search and 118 from the QGS), we conducted a multi-stage study selection process to identify the most relevant and high-quality papers for our systematic review.
@@ 2.3.1. Coarse-Grained Inclusion and Exclusion Criteria
To select relevant papers for our research questions, we defined four inclusion criteria and eight exclusion criteria (as listed in Table 2) for the coarse-grained paper selection process. Among them, In#1, Ex#1, Ex#2, and Ex#3 were automatically filtered based on the keywords, duplication status, length, and publication venue of the papers. The remaining inclusion criteria (In#2~4) and exclusion criteria (Ex#4~8) were manually applied by inspecting the topic and content of each paper. Specifically, the criteria of In#1 retained 10,302 papers whose titles and abstracts contained a pair of the identified search keywords. Subsequently, Ex#1 filtered out 797 duplicate or multi-version papers from the same authors with little difference. Next, the automated filtering criteria Ex#2 was applied to exclude short papers, tool demos, keynotes, editorials, books, theses, workshop papers, or poster papers, resulting in 6,507 papers being removed. The remaining papers were then screened based on the criteria Ex#3, which retained 622 full research papers published in the identified venues or as preprints on arXiv. The remaining inclusion and exclusion criteria (In#2~4, Ex#4~8) were then manually applied to the titles and abstracts of these 622 papers, in order to determine their relevance to the research topic. Three researchers independently applied the inclusion and exclusion criteria to the titles and abstracts. Disagreements were resolved through discussion and consensus. After this manual inspection stage, 236 papers were included for further fine-grained full-text quality assessment.
@@ 2.3.2. Fine-grained Quality Assessment
Each criterion was scored on a 3-point scale (0: not met, 1: partially met, 2: fully met). Papers with a total score of 6 or higher (out of 10) were considered as having acceptable quality. After the quality assessment, 151 papers remained in the selected set.
@@ 2.3.3. Forward and Backward Snowballing
To further expand the coverage of relevant literature, we performed forward and backward snowballing on the 151 selected papers. Forward snowballing identified papers that cited the selected papers, while backward snowballing identified papers that were referenced by the selected papers.
Here we obtained 1,072 and 2,927 papers separately during the forward and backward processes. Then we applied the same inclusion/exclusion criteria and quality assessment to the papers found through snowballing. After the initial keyword filtering and deduplication, there were 1,583 papers that remained available. Among them, 77 papers were excluded during the page number filtering step, and 1,472 papers were deleted to ensure the papers were published in the selected venues. After confirming the paper topics and assessing the paper quality, only 34 papers were ultimately retained in the snowballing process, resulting in a final set of 185 papers for data extraction and synthesis.
@@ 2.4. Statistics of Selected Papers
After conducting searches and snowballing, a total of 185 relevant research papers were ultimately obtained. The distribution of the included documents is outlined in Figure 3. As depicted in Figure 3(A), approximately 75% of the papers are from peer-reviewed venues or other sources, while the remaining 25% were published on arXiv, an open-access platform serving as a repository for scholarly articles. Among the specified peer-reviewed venues, ISSTA had the highest frequency (10%), followed by ICSE (8%) and ASE (7%). Other venues making significant contributions included FSE (4%), NDSS (3%), ACL (3%), TDSC (3%), S&P (2%), CCS (2%), USENIX Security (2%), TIFS (2%), NeurIPS (2%), EMNLP (1%), and TSE (1%). The significant portion from arXiv is unsurprising given the rapid emergence of new LLM4Security studies, with many works recently completed and potentially undergoing peer review. Despite lacking peer review in some cases, we conducted rigorous quality assessments on all collected papers to ensure the integrity of our investigation results. This approach enables us to include all high-quality and relevant publications while upholding stringent research standards.
@@
Figure 3. Overview of the selected 185 papers’ distribution.
@@ 4 The types and features of datasets used
The temporal distribution of the included papers is depicted in Figure 3(B). Since 2020, there has been a notable upward trend in the number of publications. In 2020, only 1 relevant paper was published, followed by 2 in 2021, and 11 papers in 2022. Subsequently, the field experienced explosive growth, with the total count surging to 82 papers in 2023 and a similar peak of 83 papers in 2024. This rapid growth trend signifies an increasing interest in LLM4Security research. Our review also includes 6 early papers from 2025, reflecting the ongoing momentum. We will continue to observe the developments in LLM4Security research.
@@ 3. RQ1: What types of security tasks have been facilitated by LLM-based approaches?
This section delves into the detailed examination of LLM utilization across diverse security domains. We have classified them into six primary domains, aligning with the themes of the collected papers: software and system security, network security, information and content security, hardware security, and blockchain security, totaling 185 papers. Figure 4 visually depicts the distribution of LLMs within these six domains. Additionally, Table 4 offers a comprehensive breakdown of research detailing specific security tasks addressed through LLM application.
@@
The majority of research activity in the realm of software and system security, constituting around 63% of the total research output, is attributed to the advancements made by code LLMs (Sadik et al., 2023; Zheng et al., 2024a; Zhang et al., 2024a) and the extensive applications of LLMs in software engineering (Hou et al., 2024). This emphasis underscores the significant role and impact of LLMs in software and system security, indicating a predominant focus on leveraging LLMs to automate the handling of potential security issues in programs and systems. Approximately 14% of the research focus pertains to network security tasks, highlighting the importance of LLMs in aiding traffic detection and network threat analysis. Information and content security activities represent around 12% of the research output, signaling a growing interest in employing LLMs for generating and detecting fake content. Conversely, activities in hardware security and blockchain security account for approximately 5% and 6% of the research output, respectively, suggesting that while exploration in these domains has been comparatively limited thus far, there remains research potential in utilizing LLMs to analyze hardware-level vulnerabilities and potential security risks in blockchain technology.
@@
Network Security Web fuzzing (3) 26
@@
Protocol analysis (1)
Protocol fuzzing (2)
Web vulnerability detection (1)
Software and System Security Vulnerability detection (22) 119
@@
Secure code generation (2)
ADS security verification (2)
Code obfuscation (1)
Information and Content Security Phishing and scam detection (8) 20
@@
Hardware Security Hardware vulnerability detection (2) 9
@@
Hardware IP protection (1)
Security assertions generation (1)
Blockchain Security Smart contract vulnerability detection (10) 11